xid-140091261_1Total Utility from Y Total Utility from X 24 42 56 Quantity 85 130 160 185 74 31 84 210 215

Refer to the table above. Assume that the price of good X is $2 per unit, the price of good Y is $5 per unit, and that the consumer spends a total of $44 on both goods.  What combination of X and Y will the consumer purchase in order to maximize utility?

A.

We don’t have enough information of determine his consumption bundle.

B.

He will spend more on X than on Y because X is less expensive than Y.

C.

He will purchase 4 units of Y and 4 units of X.

D.

He will purchase 2 units of Y and 2 units of X.

E.

He will purchase 6 units of Y and 7 units of X.
Answer #1

Option (E).

Utility is maximized when (MUX/PX) = (MUY/PY), where

MU = Change in Total utility (TU) / Change in Quantity (Q)

Budget constraint: 44 = 2X + 5Y

QX TUX MUX MUX/PX QY TUY MUY MUY/PY
1 24 1 85
2 42 18 9 2 130 45 9
3 56 14 7 3 160 30 6
4 66 10 5 4 185 25 5
5 74 8 4 5 200 15 3
6 80 6 3 6 210 10 2
7 84 4 2 7 215 5 1

Therefore, (MUX/PX) = (MUY/PY) is satisfied for following bundles:

(1) (MUX/PX) = (MUY/PY) = 9 when X = 2, Y = 2 and Total cost ($) = 2 x 2 + 5 x 2 = 4 + 10 = 14 < Budget

(2) (MUX/PX) = (MUY/PY) = 5 when X = 4, Y = 4 and Total cost ($) = 2 x 4 + 5 x 4 = 8 + 20 = 28 < Budget

(3) (MUX/PX) = (MUY/PY) = 3 when X = 6, Y = 5 and Total cost ($) = 2 x 6 + 5 x 5 = 12 + 25 = 37 < Budget

(4) (MUX/PX) = (MUY/PY) = 2 when X = 7, Y = 6 and Total cost ($) = 2 x 7 + 5 x 6 = 14 + 30 = 44 = Budget